OpenStudy (blaque678):

Evaluate the function at the indicated values. f(x) = 6x2 + 2x − 12 f(x + 1)= & f(−x) =

OpenStudy (tkhunny):

\(f(x) = 6x^{2} + 2x - 12\) \(f(3) = 6(3)^{2} + 2(3) - 12\) \(f(Dog) = 6(Dog)^{2} + 2(Dog) - 12\) \(f(g(x)) = 6(g(x))^{2} + 2(g(x)) - 12\) Your turn. \(f(-x) = ??\)

OpenStudy (blaque678):

what that just really confused me

OpenStudy (tkhunny):

Whatever is in those parentheses on the left, substitute for the variable on the right - no matter what it is. No confusion needed. Just do the substitution.

OpenStudy (blaque678):

the problem doesn't involve a 3 or anything with a dog or f(g(x)).That really wasn't any help at all. That just made your answer more crazy and doesn't seem like you know anything about this problem

OpenStudy (tkhunny):

It's called an example. That's how ALL such functional notation works. f(x) = 6x2 + 2x − 12 WHATEVER goes in here: f(Whatever) ALSO goes whereever it appeared in the defintion: 6(Whatever)2 + 2(Whatever) − 12 So, what do you suppose you should do with f(-x)? So, what do you suppose you should do with f(x+1)?